OpenStudy (anonymous):

What is the slope of the line passing through the points (–1, 3) and (4, –7)? a. 2 b. 3/4 c. - 4/3 d. -2

jimthompson5910 (jim_thompson5910):

slope formula \[\Large m = \frac{y_{2}-y_{1}}{x_{2}-x_{1}}\]

OpenStudy (anonymous):

(-1, 3) and (4, 7) x1, y1 x2, y2 M= y2 - y1 ----------- x2 - x1 = 7-(3) - ------- 4 - (1) = 4/5

OpenStudy (anonymous):

i am confused here

OpenStudy (anonymous):

at this part i mean

jimthompson5910 (jim_thompson5910):

In english: Step 1) subtract the y coordinates Step 2) subtract the x coordinates (same order as you did y) step 3) divide the y difference (from step 1) over the x difference (in step 2)

jimthompson5910 (jim_thompson5910):

you should have 4 - (-1) in the bottom

OpenStudy (anonymous):

yeah i do i forgot the negative here one sec

OpenStudy (anonymous):

and 4 - (-1) = 5

jimthompson5910 (jim_thompson5910):

ok where are you stuck?

OpenStudy (anonymous):

which is 4/5

OpenStudy (anonymous):

the fact that i can not simplifiy the 4/5 any lower and it is not n answer choice

jimthompson5910 (jim_thompson5910):

what are your choices?

jimthompson5910 (jim_thompson5910):

oh wait

OpenStudy (anonymous):

the ones up top

jimthompson5910 (jim_thompson5910):

you did 7-3 in the top when it should be -7-3

OpenStudy (anonymous):

OMG

jimthompson5910 (jim_thompson5910):

when you wrote `(-1, 3) and (4, 7)` you wrote the second point incorrectly so that may be why

OpenStudy (anonymous):

let me re work

OpenStudy (anonymous):

-7 - (3) ---------- 4 - (-1) = -10 ------ 5

jimthompson5910 (jim_thompson5910):

looks much better

OpenStudy (anonymous):

which is -2

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (anonymous):

haha thank you for chatching that here is a medal

jimthompson5910 (jim_thompson5910):

you're welcome